OpenStudy (anonymous):

Solve the system by graphing 3x=y+3 6x-3=2y

OpenStudy (ny,ny):

hi, i just learned systems of equations today in class. i really like them :D so, the first thing you have to do is to get these equations into y=mx+b form. then, plot the y intercept onto the graph. use the slope to find the direction of the lines. and there, you have graphed them. do you need to find a solution, too?

OpenStudy (ny,ny):

do you understand?

OpenStudy (anonymous):

No i am so confused. i need to know the points to plot.

OpenStudy (ny,ny):

ok, do you know how to put these equations in y=mx+b form?

OpenStudy (anonymous):

yea y=3x+3

OpenStudy (ny,ny):

y=-3x+3 because we subtract.

OpenStudy (ny,ny):

how about the other equation?

OpenStudy (anonymous):

2y=6x-3

OpenStudy (ny,ny):

simplify it further, you need to get the y by itself.

OpenStudy (anonymous):

divide by 2 right?

OpenStudy (ny,ny):

yes

OpenStudy (anonymous):

y=3x-3/2

OpenStudy (anonymous):

?

OpenStudy (ny,ny):

yes. good. ok so we now know the y intercept and slopes for bothe equations

OpenStudy (ny,ny):

both*

OpenStudy (ny,ny):

do you know them?

OpenStudy (anonymous):

-3=m

OpenStudy (anonymous):

would it be 3/1

OpenStudy (ny,ny):

for the second equation, the slope is 3, or 3/.

OpenStudy (ny,ny):

3/1* sorry.

OpenStudy (ny,ny):

what about the y intercepts now?

OpenStudy (anonymous):

what is the first eqution then. I thought 3/1 was the first equation?

OpenStudy (ny,ny):

the first equation is y=-3x+3. so the slope is -3. the second equation is y=3x-3/2, so the slope is 3.

OpenStudy (anonymous):

yea but the rise over run for the first is 3/1?

OpenStudy (ny,ny):

-3/1, you have to take the negative too.

OpenStudy (ny,ny):

ok?

OpenStudy (anonymous):

Yes and then the other one is -3/2?

OpenStudy (ny,ny):

the second equation is y=3x-3/2, so the slope is 3.

OpenStudy (ny,ny):

-3/2 is the y intercept.

OpenStudy (ny,ny):

ok?

OpenStudy (ny,ny):

the number next to the x is the slope. the number being added/subtracted to it is the y intercept.

OpenStudy (ny,ny):

are you getting it? i can show you everything, if that would help.

OpenStudy (anonymous):

What would the coordinate points

OpenStudy (anonymous):

Woudl you be able to help me with another problem?

OpenStudy (anonymous):

x-3y=-3 X=3y=9

OpenStudy (ny,ny):

the y intercept for the first equation is (0,3) the y intercept for the second equation is a little weird. -3/2 = -1 and 1/2. so you will plot (0,3) and (0, -1 half) always plot the y intercept first. ok? and yes I will try my best to help you with another problem :D

OpenStudy (anonymous):

Ok thank you. We sent the message at the same time. The problem is x-3y=-3 3y=x-3

OpenStudy (anonymous):

so far i have y=x/3-1

OpenStudy (ny,ny):

ok but do you have the first problem graphed?

OpenStudy (anonymous):

nvm this is the problem x-3y=-3 x+3y=9

OpenStudy (anonymous):

yes i did get it graphed

OpenStudy (ny,ny):

ok good. x-3y=-3 -3y=x-3 y=1/-3x+1 see what i did there?

OpenStudy (anonymous):

yes i did!

OpenStudy (anonymous):

I got the same answer!

OpenStudy (ny,ny):

awesome! can you write the correct equation for the second equation?

OpenStudy (ny,ny):

for x+3y=9

OpenStudy (anonymous):

y=x/3-3

OpenStudy (ny,ny):

you're close! x/3 is hard to plot because we need to know the rise, and we cant plot the rise because x is just a variable. soo... x/3 = 1/3x now that's easier to work with. that is your slope.

OpenStudy (anonymous):

-3 is my plot and 1/3 is the rise over run?

OpenStudy (ny,ny):

and... it is not -3, it is 3 because you divide 9 by 3 and that gives you 3.

OpenStudy (anonymous):

oh ok

OpenStudy (ny,ny):

got it?

OpenStudy (anonymous):

i always get my graphing wrong? I am doing it wrong somehow

OpenStudy (ny,ny):

plot 3 then move up one and to the right three points.

OpenStudy (ny,ny):

plot (0,3) i meant.

OpenStudy (anonymous):

I am confused. so what do i plot first?

OpenStudy (anonymous):

i plotted (1,0) then went up one and over -3

OpenStudy (ny,ny):